pt.A moves with velocity v which is aimed at pt.B which inturn moves rectilinearly with velocity u.initially u and v are perpendicular.how soon will they converge? please solve this using relative velocity concept.

Grade:

1 Answers

Chetan Mandayam Nayakar
12 years ago

speed at which A approaches B is constant and equal to lvl-lul, both v and u being vectors. Let the initial separation be 'r',

lvl-lul =r/t, t=r/(lvl-lul)